# Hysteresis Activation Function for Efficient Inference

### v1

@@ 2 Related Work
To increase throughput and reduce energy consumption in growing DNN models for deployment on edge devices, common methods include pruning and quantization. Pruning techniques remove less significant weights or features, requiring sophisticated re-routing in hardware to achieve actual benefits (Diao et al., 2023; Pasandi et al., 2020). Quantization, another prominent approach, reduces the precision of weights and activations to fixed precisions such as INT8 or INT4, or even binary representation (Banner et al., 2019; Rozen et al., 2022), and mixed-precision (Micikevicius et al., 2018a; Rakka et al., 2022; Kimhi et al., 2024a). This reduces the memory footprint and enables faster matrix multiplications, which are the most common operations in DNNs.
@@ 5 Conclusion
In this study, we advocate for using shifted activation threshold during training, HeLU, a novel hysteresis activation function designed to enhance the robustness of activations against the “dying ReLU” phenomenon. By leveraging the gradients of a shifted activation function, our approach maintains simplicity while improving performance across various domains. Our empirical results indicate that a relatively straightforward modification can yield significant performance gains, highlighting the potential of HeLU for universal application. This study paves the way for exploring efficient activation functions that are both computationally and environmentally sustainable. While extremely efficient, one drawback of HeLU is the search for the optimal hyper-parameter; Future research should further investigate the pre-activation statistics and find an optimal value for $\alpha$. Explore the impact on other domains such as semi supervision and segmentation Kimhi et al. (2023, 2024b) could reveal additional benefits due to the noisy training process. It is important to note that HeLU is intended to enhance efficiency rather than replace activation functions in systems where accuracy is paramount.

### v2

@@ 2 Related Work
To increase throughput and reduce energy consumption in growing DNN models for deployment on edge devices, common methods include pruning and quantization. Pruning techniques remove less significant weights or features, requiring sophisticated re-routing in hardware to achieve actual benefits (Diao et al., 2023; Pasandi et al., 2020). Quantization, another prominent approach, reduces the precision of weights and activations to fixed precisions such as INT8 or INT4, or even binary representation (Banner et al., 2019; Rozen et al., 2022), and mixed-precision (Micikevicius et al., 2018a; Rakka et al., 2022; Kimhi et al., 2024b). This reduces the memory footprint and enables faster matrix multiplications, which are the most common operations in DNNs.
@@ 5 Conclusion
In this study, we advocate for using shifted activation threshold during training, HeLU, a novel hysteresis activation function designed to enhance the robustness of activations against the “dying ReLU” phenomenon. By leveraging the gradients of a shifted activation function, our approach maintains simplicity while improving performance across various domains. Our empirical results indicate that a relatively straightforward modification can yield significant performance gains, highlighting the potential of HeLU for universal application. This study paves the way for exploring efficient activation functions that are both computationally and environmentally sustainable. While extremely efficient, one drawback of HeLU is the search for the optimal hyper-parameter; Future research should further investigate the pre-activation statistics and find an optimal value for $\alpha$. Explore the impact on other domains such as segmentation, specifically under reduced supervision Kimhi et al. (2024a, 2025) or noisy labeled Grad et al. (2024), could reveal additional benefits due to the noisy training process. It is important to note that HeLU is intended to enhance efficiency rather than replace activation functions in systems where accuracy is paramount.